# Hallucination Detection in LLMs Using Spectral Features of Attention Maps

### v1

@@ Appendix B Details of QA datasets
Figure 8: Token count histograms for datasets used in the experiments. Number of tokens determined separately for $question$ (left-hand-side plots) and gold $answer$ (right-hand-side plots) of each example in the datasets with meta-llama/Llama-3.1-8B-Instruct tokeniser (whenever multiple possible answers occurred, they were flattened).
Appendix C Hallucination dataset sizes
@@
In Figure 9, we show the number of examples for each label determined with the llm-as-judge heuristic. It is worth noting that different generation configurations result in different splits, as LLMs might produce different answers. All examples classified as $Rejected$ were discarded from further experiments. We can observe that in most cases, datasets are imbalanced, underrepresenting non-hallucinated examples. Only for TriviaQA, there is an approximately balanced number of examples or even more non-hallucinated ones, depending on the configuration used. We split each dataset into 80% training examples and 20% test examples. Splits were stratified according to hallucination labels.
@@
Figure 9: Number of examples per each label in generated datasets ($Hallucination$ number of hallucinated examples, $Non{-}Hallucination$ number of truthful examples, $Rejected$ number of examples unable to evaluate).
Appendix D Extended results
D.1 Extended method comparison
@@
In Table 4, we present the extended results from Table 1 in the main part of this paper. These results cover probes trained with both all-layers and per-layer variants across all models, as well as lower temperature ($temp\in\{0.1,1.0\}$). In all cases, the all-layers variant outperforms the per-layer variant, suggesting that hallucination-related information is distributed across multiple layers. Additionally, we observe a smaller generalisation gap (measured as the difference between test and training performance) for the $\operatorname{LapEigvals}$ method, indicating more robust features present in the Laplacian eigenvalues. Finally, as demonstrated in Section 6, increasing the temperature during answer generation improves probe performance, which is also evident in Table 4, where probes trained on answers generated with $temp=1.0$ consistently outperform those trained on data generated with $temp=0.1$.
@@
Table 4: Performance comparison of methods evaluated in this work on an extended set of configurations. We marked results for $\operatorname{AttentionScore}$ in gray as it is unsupervised approach, not directly comparable to the other ones. In bold, we highlight the best performance on the test split of data, individually for each dataset, LLM, and temperature.
@@
CoQA HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA CoQA HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.1 $\operatorname{AttentionScore}$ 0.509 0.667 0.607 0.556 0.567 0.563 0.541 0.653 0.631 0.575 0.571 0.650
Llama3.1-8B 0.1 $\operatorname{AttentionScore}$ 0.494 0.614 0.568 0.522 0.522 0.489 0.504 0.587 0.558 0.521 0.511 0.537
Llama3.1-8B 0.1 $\operatorname{AttnLogDet}$ 0.574 0.776 0.702 0.688 0.739 0.709 0.606 0.770 0.713 0.708 0.741 0.777
Llama3.1-8B 0.1 $\operatorname{AttnLogDet}$ 0.843 0.884 0.851 0.839 0.861 0.913 0.770 0.837 0.768 0.758 0.827 0.820
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 0.764 0.828 0.713 0.742 0.793 0.680 0.729 0.799 0.728 0.749 0.773 0.790
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 0.861 0.895 0.878 0.858 0.867 0.979 0.776 0.838 0.755 0.781 0.822 0.819
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 0.758 0.817 0.698 0.707 0.781 0.708 0.757 0.793 0.711 0.733 0.780 0.764
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 0.869 0.901 0.864 0.855 0.896 0.903 0.836 0.867 0.793 0.782 0.872 0.822
Llama3.1-8B 1.0 $\operatorname{AttentionScore}$ 0.514 0.640 0.607 0.558 0.578 0.533 0.525 0.642 0.607 0.572 0.602 0.629
Llama3.1-8B 1.0 $\operatorname{AttentionScore}$ 0.507 0.602 0.580 0.534 0.535 0.546 0.493 0.589 0.556 0.538 0.532 0.541
Llama3.1-8B 1.0 $\operatorname{AttnLogDet}$ 0.596 0.755 0.704 0.697 0.750 0.757 0.597 0.763 0.757 0.686 0.754 0.771
Llama3.1-8B 1.0 $\operatorname{AttnLogDet}$ 0.848 0.882 0.856 0.846 0.867 0.930 0.769 0.827 0.793 0.748 0.842 0.814
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 0.762 0.820 0.758 0.754 0.800 0.796 0.723 0.784 0.732 0.728 0.796 0.770
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 0.867 0.889 0.873 0.867 0.876 0.972 0.782 0.819 0.790 0.768 0.843 0.833
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 0.760 0.803 0.732 0.722 0.795 0.751 0.743 0.789 0.725 0.724 0.794 0.764
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 0.879 0.896 0.866 0.857 0.901 0.918 0.830 0.874 0.827 0.791 0.889 0.829
Llama3.2-3B 0.1 $\operatorname{AttentionScore}$ 0.526 0.697 0.592 0.570 0.570 0.569 0.547 0.714 0.643 0.582 0.551 0.564
Llama3.2-3B 0.1 $\operatorname{AttentionScore}$ 0.506 0.635 0.523 0.515 0.534 0.473 0.519 0.644 0.573 0.561 0.510 0.489
Llama3.2-3B 0.1 $\operatorname{AttnLogDet}$ 0.573 0.762 0.692 0.682 0.719 0.725 0.579 0.774 0.735 0.698 0.711 0.674
Llama3.2-3B 0.1 $\operatorname{AttnLogDet}$ 0.782 0.868 0.845 0.827 0.824 0.918 0.695 0.843 0.763 0.749 0.796 0.678
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 0.675 0.782 0.750 0.725 0.755 0.727 0.626 0.792 0.734 0.695 0.724 0.720
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 0.814 0.873 0.872 0.852 0.842 0.963 0.723 0.844 0.772 0.744 0.788 0.688
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 0.681 0.774 0.733 0.708 0.733 0.722 0.676 0.781 0.736 0.697 0.732 0.690
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 0.831 0.875 0.837 0.832 0.852 0.895 0.801 0.857 0.779 0.736 0.826 0.743
Llama3.2-3B 1.0 $\operatorname{AttentionScore}$ 0.532 0.668 0.588 0.578 0.553 0.555 0.557 0.637 0.592 0.593 0.558 0.675
Llama3.2-3B 1.0 $\operatorname{AttentionScore}$ 0.512 0.606 0.554 0.529 0.517 0.484 0.509 0.588 0.546 0.530 0.515 0.581
Llama3.2-3B 1.0 $\operatorname{AttnLogDet}$ 0.578 0.738 0.677 0.720 0.716 0.739 0.597 0.724 0.678 0.707 0.711 0.742
Llama3.2-3B 1.0 $\operatorname{AttnLogDet}$ 0.784 0.869 0.816 0.839 0.831 0.924 0.700 0.801 0.690 0.734 0.789 0.795
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 0.642 0.777 0.716 0.747 0.763 0.735 0.641 0.756 0.696 0.703 0.746 0.748
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 0.819 0.878 0.847 0.876 0.847 0.978 0.724 0.819 0.694 0.749 0.804 0.723
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 0.695 0.764 0.683 0.719 0.727 0.682 0.715 0.754 0.671 0.711 0.738 0.767
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 0.842 0.885 0.803 0.850 0.863 0.911 0.812 0.828 0.693 0.757 0.832 0.787
Phi3.5 0.1 $\operatorname{AttentionScore}$ 0.517 0.559 0.565 0.606 0.625 0.601 0.528 0.551 0.637 0.621 0.628 0.637
Phi3.5 0.1 $\operatorname{AttentionScore}$ 0.499 0.538 0.532 0.473 0.539 0.522 0.505 0.511 0.578 0.458 0.534 0.554
Phi3.5 0.1 $\operatorname{AttnLogDet}$ 0.583 0.732 0.741 0.711 0.757 0.720 0.585 0.726 0.785 0.726 0.772 0.765
Phi3.5 0.1 $\operatorname{AttnLogDet}$ 0.845 0.863 0.905 0.852 0.875 0.981 0.723 0.802 0.802 0.759 0.842 0.716
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 0.760 0.781 0.793 0.745 0.802 0.854 0.678 0.764 0.790 0.747 0.791 0.774
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 0.862 0.867 0.904 0.861 0.881 0.999 0.728 0.802 0.787 0.740 0.838 0.761
Phi3.5 0.1 $\operatorname{LapEigvals}$ 0.734 0.758 0.737 0.704 0.775 0.759 0.716 0.757 0.761 0.732 0.768 0.741
Phi3.5 0.1 $\operatorname{LapEigvals}$ 0.856 0.860 0.897 0.841 0.884 0.965 0.810 0.819 0.815 0.791 0.858 0.717
Phi3.5 1.0 $\operatorname{AttentionScore}$ 0.499 0.567 0.615 0.626 0.637 0.618 0.533 0.581 0.630 0.645 0.642 0.626
Phi3.5 1.0 $\operatorname{AttentionScore}$ 0.489 0.540 0.566 0.469 0.553 0.541 0.520 0.541 0.594 0.504 0.540 0.554
Phi3.5 1.0 $\operatorname{AttnLogDet}$ 0.587 0.733 0.773 0.722 0.766 0.753 0.557 0.762 0.784 0.736 0.772 0.763
Phi3.5 1.0 $\operatorname{AttnLogDet}$ 0.842 0.868 0.921 0.859 0.879 0.971 0.745 0.818 0.815 0.769 0.848 0.755
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 0.755 0.794 0.820 0.790 0.809 0.864 0.710 0.795 0.787 0.752 0.799 0.747
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 0.858 0.871 0.924 0.876 0.887 0.998 0.771 0.829 0.798 0.782 0.850 0.802
Phi3.5 1.0 $\operatorname{LapEigvals}$ 0.733 0.755 0.755 0.718 0.779 0.713 0.723 0.769 0.755 0.732 0.792 0.732
Phi3.5 1.0 $\operatorname{LapEigvals}$ 0.856 0.863 0.911 0.849 0.889 0.961 0.821 0.836 0.826 0.795 0.872 0.777
D.2 Best found hyperparameters

### v2

@@ Appendix D Details of QA datasets
Figure 8: Token count histograms for the datasets used in our experiments. Token counts were computed separately for each example’s questionquestion (left) and gold answeranswer (right) using the meta-llama/Llama-3.1-8B-Instruct tokenizer. In cases with multiple answers, they were flattened into one.
Appendix E Hallucination dataset sizes
@@
Figure 9 shows the number of examples per label, determined using exact match for GSM8K and the llm-as-judge heuristic for the other datasets. It is worth noting that different generation configurations result in different splits, as LLMs might produce different answers. All examples classified as $Rejected$ were discarded from the hallucination probe training and evaluation. We observe that most datasets are imbalanced, typically underrepresenting non-hallucinated examples, with the exception of TriviaQA and GSM8K. We split each dataset into 80% training examples and 20% test examples. Splits were stratified according to hallucination labels.
@@
Figure 9: Number of examples per each label in generated datasets (HallucinationHallucination - number of hallucinated examples, NonHallucinationNon{-}Hallucination - number of truthful examples, RejectedRejected - number of examples unable to evaluate).
Appendix F LLM-as-Judge agreement
To ensure the high quality of labels generated using the llm-as-judge approach, we complemented manual evaluation of random examples with a second judge LLM and measured agreement between the models. We assume that higher agreement among LLMs indicates better label quality. The reduced performance of $\operatorname{LapEigvals}$ on TriviaQA may be attributed to the lower agreement, as well as the dataset’s size and class imbalance discussed earlier.
Table 4: Agreement between LLM judges labeling hallucinations (gpt-4o-mini, gpt-4.1), measured with Cohen’s Kappa.
Dataset Cohen’s Kappa
CoQA 0.876
HaluevalQA 0.946
NQOpen 0.883
SquadV2 0.854
TriviaQA 0.939
TruthfulQA 0.714
Appendix G Extended results
G.1 Precision and Recall analysis
To provide insights relevant for potential practical usage, we analyze the Precision and Recall of our method. While it has not yet been fully evaluated in production settings, this analysis illustrates the trade-offs between these metrics and informs how the method might behave in real-world applications. Metrics were computed using the default threshold of 0.5, as reported in Table 5. Although trade-off patterns vary across datasets, they are consistent across all evaluated LLMs. Specifically, we observe higher recall on CoQA, GSM8K, and TriviaQA, whereas HaluEvalQA, NQ-Open, SQuADv2, and TruthfulQA exhibit higher precision. These insights can guide threshold adjustments to balance precision and recall for different production scenarios.
Table 5: Precision and Recall values for the $\operatorname{LapEigvals}$ method, complementary to AUROC presented in Table 1. Values are presented as Precision / Recall for each dataset and model combination.
LLM CoQA GSM8K HaluEvalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.583 / 0.710 0.644 / 0.729 0.895 / 0.785 0.859 / 0.740 0.896 / 0.720 0.719 / 0.812 0.872 / 0.781
Llama3.2-3B 0.679 / 0.728 0.718 / 0.699 0.912 / 0.788 0.894 / 0.662 0.924 / 0.720 0.787 / 0.729 0.910 / 0.746
Phi3.5 0.560 / 0.703 0.600 / 0.739 0.899 / 0.768 0.910 / 0.785 0.906 / 0.731 0.787 / 0.785 0.829 / 0.798
Mistral-Nemo 0.646 / 0.714 0.594 / 0.809 0.873 / 0.760 0.875 / 0.751 0.920 / 0.756 0.707 / 0.769 0.892 / 0.825
Mistral-Small-24B 0.610 / 0.779 0.561 / 0.852 0.811 / 0.801 0.700 / 0.750 0.784 / 0.789 0.575 / 0.787 0.679 / 0.655
G.2 Extended method comparison
@@
In Tables 6 and 7, we present the extended results corresponding to those summarized in Table 1 in the main part of this paper. The extended results cover probes trained with both all-layers and per-layer variants across all models, as well as lower temperature ($temp\in\{0.1,1.0\}$). In almost all cases, the all-layers variant outperforms the per-layer variant, suggesting that hallucination-related information is distributed across multiple layers. Additionally, we observe a smaller generalization gap (measured as the difference between test and training performance) for the $\operatorname{LapEigvals}$ method, indicating more robust features present in the Laplacian eigenvalues. Finally, as demonstrated in Section 6, increasing the temperature during answer generation improves probe performance, which is also evident in Table 6, where probes trained on answers generated with temp=1.0temp{=}1.0 consistently outperform those trained on data generated with temp=0.1temp{=}0.1.
@@
Table 6: (Part I) Performance comparison of methods on an extended set of configurations. We mark results for $\operatorname{AttentionScore}$ in gray as it is an unsupervised approach, not directly comparable to the others. In bold, we highlight the best performance on the test split of data, individually for each dataset, LLM, and temperature.
@@
CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B 0.1 $\operatorname{AttentionScore}$ 0.509 0.683 0.667 0.607 0.556 0.567 0.563 0.541 0.764 0.653 0.631 0.575 0.571 0.650
Llama3.1-8B 0.1 $\operatorname{AttentionScore}$ 0.494 0.677 0.614 0.568 0.522 0.522 0.489 0.504 0.708 0.587 0.558 0.521 0.511 0.537
Llama3.1-8B 0.1 $\operatorname{AttnLogDet}$ 0.574 0.810 0.776 0.702 0.688 0.739 0.709 0.606 0.840 0.770 0.713 0.708 0.741 0.777
Llama3.1-8B 0.1 $\operatorname{AttnLogDet}$ 0.843 0.977 0.884 0.851 0.839 0.861 0.913 0.770 0.833 0.837 0.768 0.758 0.827 0.820
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 0.764 0.879 0.828 0.713 0.742 0.793 0.680 0.729 0.798 0.799 0.728 0.749 0.773 0.790
Llama3.1-8B 0.1 $\operatorname{AttnEigvals}$ 0.861 0.992 0.895 0.878 0.858 0.867 0.979 0.776 0.841 0.838 0.755 0.781 0.822 0.819
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 0.758 0.777 0.817 0.698 0.707 0.781 0.708 0.757 0.844 0.793 0.711 0.733 0.780 0.764
Llama3.1-8B 0.1 $\operatorname{LapEigvals}$ 0.869 0.928 0.901 0.864 0.855 0.896 0.903 0.836 0.887 0.867 0.793 0.782 0.872 0.822
Llama3.1-8B 1.0 $\operatorname{AttentionScore}$ 0.514 0.705 0.640 0.607 0.558 0.578 0.533 0.525 0.731 0.642 0.607 0.572 0.602 0.629
Llama3.1-8B 1.0 $\operatorname{AttentionScore}$ 0.507 0.710 0.602 0.580 0.534 0.535 0.546 0.493 0.720 0.589 0.556 0.538 0.532 0.541
Llama3.1-8B 1.0 $\operatorname{AttnLogDet}$ 0.596 0.791 0.755 0.704 0.697 0.750 0.757 0.597 0.828 0.763 0.757 0.686 0.754 0.771
Llama3.1-8B 1.0 $\operatorname{AttnLogDet}$ 0.848 0.973 0.882 0.856 0.846 0.867 0.930 0.769 0.826 0.827 0.793 0.748 0.842 0.814
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 0.762 0.864 0.820 0.758 0.754 0.800 0.796 0.723 0.812 0.784 0.732 0.728 0.796 0.770
Llama3.1-8B 1.0 $\operatorname{AttnEigvals}$ 0.867 0.995 0.889 0.873 0.867 0.876 0.972 0.782 0.838 0.819 0.790 0.768 0.843 0.833
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 0.760 0.873 0.803 0.732 0.722 0.795 0.751 0.743 0.833 0.789 0.725 0.724 0.794 0.764
Llama3.1-8B 1.0 $\operatorname{LapEigvals}$ 0.879 0.936 0.896 0.866 0.857 0.901 0.918 0.830 0.872 0.874 0.827 0.791 0.889 0.829
Llama3.2-3B 0.1 $\operatorname{AttentionScore}$ 0.526 0.662 0.697 0.592 0.570 0.570 0.569 0.547 0.640 0.714 0.643 0.582 0.551 0.564
Llama3.2-3B 0.1 $\operatorname{AttentionScore}$ 0.506 0.638 0.635 0.523 0.515 0.534 0.473 0.519 0.609 0.644 0.573 0.561 0.510 0.489
Llama3.2-3B 0.1 $\operatorname{AttnLogDet}$ 0.573 0.774 0.762 0.692 0.682 0.719 0.725 0.579 0.794 0.774 0.735 0.698 0.711 0.674
Llama3.2-3B 0.1 $\operatorname{AttnLogDet}$ 0.782 0.946 0.868 0.845 0.827 0.824 0.918 0.695 0.841 0.843 0.763 0.749 0.796 0.678
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 0.675 0.784 0.782 0.750 0.725 0.755 0.727 0.626 0.761 0.792 0.734 0.695 0.724 0.720
Llama3.2-3B 0.1 $\operatorname{AttnEigvals}$ 0.814 0.977 0.873 0.872 0.852 0.842 0.963 0.723 0.808 0.844 0.772 0.744 0.788 0.688
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 0.681 0.763 0.774 0.733 0.708 0.733 0.722 0.676 0.835 0.781 0.736 0.697 0.732 0.690
Llama3.2-3B 0.1 $\operatorname{LapEigvals}$ 0.831 0.889 0.875 0.837 0.832 0.852 0.895 0.801 0.852 0.857 0.779 0.736 0.826 0.743
Llama3.2-3B 1.0 $\operatorname{AttentionScore}$ 0.532 0.674 0.668 0.588 0.578 0.553 0.555 0.557 0.753 0.637 0.592 0.593 0.558 0.675
Llama3.2-3B 1.0 $\operatorname{AttentionScore}$ 0.512 0.648 0.606 0.554 0.529 0.517 0.484 0.509 0.717 0.588 0.546 0.530 0.515 0.581
Llama3.2-3B 1.0 $\operatorname{AttnLogDet}$ 0.578 0.807 0.738 0.677 0.720 0.716 0.739 0.597 0.816 0.724 0.678 0.707 0.711 0.742
Llama3.2-3B 1.0 $\operatorname{AttnLogDet}$ 0.784 0.951 0.869 0.816 0.839 0.831 0.924 0.700 0.851 0.801 0.690 0.734 0.789 0.795
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 0.642 0.807 0.777 0.716 0.747 0.763 0.735 0.641 0.817 0.756 0.696 0.703 0.746 0.748
Llama3.2-3B 1.0 $\operatorname{AttnEigvals}$ 0.819 0.973 0.878 0.847 0.876 0.847 0.978 0.724 0.768 0.819 0.694 0.749 0.804 0.723
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 0.695 0.781 0.764 0.683 0.719 0.727 0.682 0.715 0.815 0.754 0.671 0.711 0.738 0.767
Llama3.2-3B 1.0 $\operatorname{LapEigvals}$ 0.842 0.894 0.885 0.803 0.850 0.863 0.911 0.812 0.870 0.828 0.693 0.757 0.832 0.787
Phi3.5 0.1 $\operatorname{AttentionScore}$ 0.517 0.723 0.559 0.565 0.606 0.625 0.601 0.528 0.682 0.551 0.637 0.621 0.628 0.637
Phi3.5 0.1 $\operatorname{AttentionScore}$ 0.499 0.632 0.538 0.532 0.473 0.539 0.522 0.505 0.605 0.511 0.578 0.458 0.534 0.554
Phi3.5 0.1 $\operatorname{AttnLogDet}$ 0.583 0.805 0.732 0.741 0.711 0.757 0.720 0.585 0.749 0.726 0.785 0.726 0.772 0.765
Phi3.5 0.1 $\operatorname{AttnLogDet}$ 0.845 0.995 0.863 0.905 0.852 0.875 0.981 0.723 0.752 0.802 0.802 0.759 0.842 0.716
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 0.760 0.882 0.781 0.793 0.745 0.802 0.854 0.678 0.764 0.764 0.790 0.747 0.791 0.774
Phi3.5 0.1 $\operatorname{AttnEigvals}$ 0.862 1.000 0.867 0.904 0.861 0.881 0.999 0.728 0.732 0.802 0.787 0.740 0.838 0.761
Phi3.5 0.1 $\operatorname{LapEigvals}$ 0.734 0.713 0.758 0.737 0.704 0.775 0.759 0.716 0.753 0.757 0.761 0.732 0.768 0.741
Phi3.5 0.1 $\operatorname{LapEigvals}$ 0.856 0.946 0.860 0.897 0.841 0.884 0.965 0.810 0.785 0.819 0.815 0.791 0.858 0.717
Phi3.5 1.0 $\operatorname{AttentionScore}$ 0.499 0.699 0.567 0.615 0.626 0.637 0.618 0.533 0.722 0.581 0.630 0.645 0.642 0.626
Phi3.5 1.0 $\operatorname{AttentionScore}$ 0.489 0.640 0.540 0.566 0.469 0.553 0.541 0.520 0.666 0.541 0.594 0.504 0.540 0.554
Phi3.5 1.0 $\operatorname{AttnLogDet}$ 0.587 0.831 0.733 0.773 0.722 0.766 0.753 0.557 0.842 0.762 0.784 0.736 0.772 0.763
Phi3.5 1.0 $\operatorname{AttnLogDet}$ 0.842 0.993 0.868 0.921 0.859 0.879 0.971 0.745 0.842 0.818 0.815 0.769 0.848 0.755
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 0.755 0.852 0.794 0.820 0.790 0.809 0.864 0.710 0.809 0.795 0.787 0.752 0.799 0.747
Phi3.5 1.0 $\operatorname{AttnEigvals}$ 0.858 1.000 0.871 0.924 0.876 0.887 0.998 0.771 0.794 0.829 0.798 0.782 0.850 0.802
Phi3.5 1.0 $\operatorname{LapEigvals}$ 0.733 0.771 0.755 0.755 0.718 0.779 0.713 0.723 0.816 0.769 0.755 0.732 0.792 0.732
Phi3.5 1.0 $\operatorname{LapEigvals}$ 0.856 0.937 0.863 0.911 0.849 0.889 0.961 0.821 0.885 0.836 0.826 0.795 0.872 0.777
Table 7: (Part II) Performance comparison of methods on an extended set of configurations. We mark results for $\operatorname{AttentionScore}$ in gray as it is an unsupervised approach, not directly comparable to the others. In bold, we highlight the best performance on the test split of data, individually for each dataset, LLM, and temperature.
LLM Temp Feature all-layers per-layer Train AUROC Test AUROC
CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Mistral-Nemo 0.1 $\operatorname{AttentionScore}$ 0.504 0.727 0.574 0.591 0.509 0.550 0.546 0.515 0.697 0.559 0.587 0.527 0.545 0.681
Mistral-Nemo 0.1 $\operatorname{AttentionScore}$ 0.508 0.707 0.536 0.537 0.507 0.520 0.535 0.484 0.667 0.523 0.533 0.495 0.505 0.631
Mistral-Nemo 0.1 $\operatorname{AttnLogDet}$ 0.584 0.801 0.716 0.702 0.675 0.689 0.744 0.583 0.807 0.723 0.688 0.668 0.722 0.731
Mistral-Nemo 0.1 $\operatorname{AttnLogDet}$ 0.828 0.993 0.842 0.861 0.858 0.854 0.963 0.734 0.820 0.786 0.752 0.709 0.822 0.776
Mistral-Nemo 0.1 $\operatorname{AttnEigvals}$ 0.708 0.865 0.751 0.749 0.749 0.747 0.797 0.672 0.795 0.740 0.701 0.704 0.738 0.717
Mistral-Nemo 0.1 $\operatorname{AttnEigvals}$ 0.845 1.000 0.842 0.878 0.864 0.859 0.996 0.768 0.771 0.789 0.743 0.716 0.809 0.752
Mistral-Nemo 0.1 $\operatorname{LapEigvals}$ 0.763 0.777 0.772 0.732 0.723 0.781 0.725 0.759 0.751 0.760 0.697 0.696 0.769 0.710
Mistral-Nemo 0.1 $\operatorname{LapEigvals}$ 0.868 0.969 0.862 0.875 0.869 0.886 0.977 0.823 0.805 0.821 0.755 0.767 0.858 0.737
Mistral-Nemo 1.0 $\operatorname{AttentionScore}$ 0.502 0.656 0.586 0.606 0.546 0.553 0.570 0.525 0.670 0.587 0.588 0.564 0.570 0.632
Mistral-Nemo 1.0 $\operatorname{AttentionScore}$ 0.493 0.675 0.541 0.552 0.503 0.521 0.531 0.493 0.630 0.531 0.529 0.510 0.532 0.494
Mistral-Nemo 1.0 $\operatorname{AttnLogDet}$ 0.591 0.790 0.723 0.716 0.717 0.717 0.741 0.581 0.782 0.730 0.703 0.711 0.707 0.801
Mistral-Nemo 1.0 $\operatorname{AttnLogDet}$ 0.829 0.994 0.851 0.870 0.860 0.857 0.963 0.728 0.856 0.798 0.769 0.772 0.812 0.852
Mistral-Nemo 1.0 $\operatorname{AttnEigvals}$ 0.704 0.845 0.762 0.742 0.757 0.752 0.806 0.670 0.781 0.749 0.742 0.719 0.737 0.804
Mistral-Nemo 1.0 $\operatorname{AttnEigvals}$ 0.844 1.000 0.851 0.893 0.864 0.862 0.996 0.778 0.842 0.781 0.761 0.758 0.821 0.802
Mistral-Nemo 1.0 $\operatorname{LapEigvals}$ 0.765 0.820 0.790 0.749 0.740 0.804 0.779 0.738 0.808 0.763 0.708 0.723 0.785 0.818
Mistral-Nemo 1.0 $\operatorname{LapEigvals}$ 0.876 0.965 0.877 0.884 0.881 0.901 0.978 0.835 0.890 0.833 0.795 0.812 0.865 0.828
Mistral-Small-24B 0.1 $\operatorname{AttentionScore}$ 0.520 0.759 0.538 0.517 0.577 0.535 0.571 0.525 0.685 0.552 0.592 0.625 0.533 0.724
Mistral-Small-24B 0.1 $\operatorname{AttentionScore}$ 0.520 0.668 0.472 0.449 0.510 0.449 0.491 0.493 0.578 0.493 0.467 0.556 0.461 0.645
Mistral-Small-24B 0.1 $\operatorname{AttnLogDet}$ 0.585 0.834 0.674 0.659 0.724 0.685 0.698 0.586 0.809 0.684 0.695 0.752 0.682 0.721
Mistral-Small-24B 0.1 $\operatorname{AttnLogDet}$ 0.851 0.990 0.817 0.799 0.820 0.861 0.898 0.762 0.896 0.760 0.725 0.763 0.778 0.767
Mistral-Small-24B 0.1 $\operatorname{AttnEigvals}$ 0.734 0.863 0.722 0.667 0.745 0.757 0.732 0.720 0.837 0.707 0.697 0.773 0.758 0.765
Mistral-Small-24B 0.1 $\operatorname{AttnEigvals}$ 0.872 0.999 0.873 0.923 0.903 0.899 0.993 0.793 0.896 0.771 0.731 0.803 0.809 0.796
Mistral-Small-24B 0.1 $\operatorname{LapEigvals}$ 0.802 0.781 0.720 0.646 0.714 0.742 0.694 0.800 0.850 0.719 0.674 0.784 0.757 0.827
Mistral-Small-24B 0.1 $\operatorname{LapEigvals}$ 0.887 0.985 0.870 0.901 0.887 0.905 0.979 0.852 0.881 0.808 0.722 0.821 0.831 0.757
Mistral-Small-24B 1.0 $\operatorname{AttentionScore}$ 0.511 0.706 0.555 0.582 0.561 0.562 0.542 0.535 0.713 0.566 0.576 0.567 0.574 0.606
Mistral-Small-24B 1.0 $\operatorname{AttentionScore}$ 0.497 0.595 0.503 0.463 0.519 0.451 0.493 0.516 0.576 0.504 0.462 0.455 0.463 0.451
Mistral-Small-24B 1.0 $\operatorname{AttnLogDet}$ 0.591 0.824 0.727 0.710 0.732 0.720 0.677 0.600 0.869 0.771 0.714 0.726 0.734 0.687
Mistral-Small-24B 1.0 $\operatorname{AttnLogDet}$ 0.850 0.989 0.847 0.827 0.856 0.853 0.877 0.766 0.853 0.842 0.747 0.753 0.833 0.735
Mistral-Small-24B 1.0 $\operatorname{AttnEigvals}$ 0.757 0.920 0.743 0.728 0.764 0.779 0.741 0.723 0.868 0.780 0.733 0.734 0.780 0.718
Mistral-Small-24B 1.0 $\operatorname{AttnEigvals}$ 0.877 1.000 0.878 0.923 0.911 0.895 0.997 0.805 0.846 0.848 0.751 0.760 0.844 0.765
Mistral-Small-24B 1.0 $\operatorname{LapEigvals}$ 0.814 0.860 0.762 0.733 0.790 0.766 0.703 0.805 0.897 0.790 0.712 0.781 0.779 0.725
Mistral-Small-24B 1.0 $\operatorname{LapEigvals}$ 0.895 0.980 0.890 0.898 0.910 0.907 0.965 0.861 0.925 0.882 0.791 0.820 0.876 0.748
G.3 Best found hyperparameters